The fire is now at one hundred percent containment and is in patrol status. Smoldering areas within the interior may continue to put up smoke. The fire will be determined completely controlled after a season-ending rain event occurs. Dozer and hand line repair have concluded in the east and west zones. Ninety eight percent of the suppression repair work is now complete with 374 out of 381 miles repaired. The seven miles of remaining road grading work and culvert repairs on the North Coast Ridge Road will be completed after weather improves. The South Central Sierra Interagency Incident Management Team transitioned command back to Los Padres National Forest Type 3 Team at 1800 hours Tuesday October 25th. The Type 3 team has continued to complete suppression repair priorities ahead of significant rain event and demobilized all remaining equipment and crews. Transfer of command to the Type 4 Incident Commander on the Monterey Ranger District will occur on Friday October 28th, 2016 @ 18:00. The Monterey Ranger District of Los Padres National Forest remains closed to public access. The only recreation sites that remain open are Plaskett Creek and Kirk Creek Campgrounds, Pfeiffer Beach, Sand Dollar Beach, and Willow Creek Day Use areas. All trails are closed within the Ventana and Silver Peak Wilderness areas.
The Burned Area Emergency Response (BAER) assessment has been completed and implementation is on-going. For more information on the Soberanes Fire BAER Team activities, visit their InciWeb page at Soberanes BAER Inciweb.
|Current as of|
|Date of Origin||Friday July 22nd, 2016 approx. 08:45 AM|
|Location||Soberanes Creek, Garrapata State Park, Palo Colorado/Big Sur, & Ventana Wilderness.|
|Incident Commander||Hubert Chip Laugharn and Brian Sexton, Los Padres Type 4 Organization|
|Incident Description||Wildland Fire|
|Percent of Perimeter Contained||100%|
Chaparral, tall grass and timber
As of October 28th, 2016 at 11:00 a.m., the ICP was relocated to the Monterey Ranger District Office in King City.
Current weather systems have delayed suppression repair operations. As weather permits, those operations will continue. The fire is in patrol status to hold it to its current footprint. Implementation of suppression repair with the resource advisor's direction continues.
|Projected Incident Activity|
No significant fire growth is expected.
Ninety nine percent of known suppression repair work has been completed. The remaining suppression repair work will continue as conditions allow.
Currently, rain is predicted for the entire fire area today through the weekend. Another system, bringing rain to the fire area is predicted for Monday night into Tuesday.